# ChipletPart: Cost-Aware Partitioning for 2.5D Systems

### v2

@@ Abstract.
journal: TODAES
@@ 1. Introduction
The integration of multiple chips on an interposer has become a favorable approach to reduce the cost of building large systems (IntelPV, ) (Sangiovanni-VincentelliLZZ23, ). In a 2.5D system, a design is decomposed into multiple smaller chiplets, which are packaged together on the same substrate. This splitting can have significant benefits for yield and can enable designs using heterogeneous process technology nodes, going beyond what is possible in a monolithic design. On the other hand, inter-die communication exhibits higher area and power overhead compared to intra-die communication. Due to this and other factors, the potential cost benefits of disaggregation are not always realized (GraeningPG23, ). Therefore, it is important to intelligently partition a design into constituent chiplets and assign a manufacturing technology to each of the chiplets so as to optimize manufacturing costs.
@@
The natural partitioning granularity for 2.5D integration is at the block-level. Splitting individual IP blocks into multiple chiplets is problematic for design and test methodology reasons as well as for IP reuse. Additionally, splitting an IP block will often have a significant impact on performance and IO count. Due to these considerations, block-level chiplet partitioning (Figure 1) is the focus of this paper. The cost-aware partitioning problem for chiplet systems is fundamentally different from the well-studied (KarypisAKS99, ) (bustany2022specpart, ) (BustanyGKKPW23, ) netlist min-cut partitioning. The problem size is smaller (usually a few hundred blocks and a few tens of chiplets at most) and the underlying objective function can be complex. Furthermore, inter-chiplet IO reach limitations can render some partitions infeasible, necessitating floorplan-awareness in the partitioner. This is further complicated by the need to select a manufacturing technology for each of the chiplets.
@@
Figure 1. Generic example of block-level chiplet partitioning. Left: an IP block-level netlist (different types of IP blocks shown in different colors). Right: an integrated 2.5D chiplet system reflecting netlist partitioning with technology node assignment shown by different colors.
[…]
@@
We propose ChipletPart, the first open-source, unified framework for chiplet partitioning that combines cost-driven multi-way partitioning, heterogeneous technology assignment and IO reach-aware floorplanning in a single, automated flow. ChipletPart presents a novel adaptation of classical optimizers such as genetic algorithms and simulated annealing to generate floorplan-feasible, cost-optimized and technology-assigned 2.5D chiplet systems. ChipletPart is intended for design space exploration before physical implementation.  [^1] ChipletPart considers a given architecture and generates partitioned chiplets before physical implementation. We do not perform architecture exploration in this work. We envision ChipletPart as a tool to assist human designers by generating strong initial solutions, which can then be refined using domain expertise. Our key contributions include:
@@
  • Integrated chiplet partitioning with floorplan feasibility: We present the first unified framework that optimizes system partitioning while guaranteeing feasibility under IO reach constraints. In contrast to standard min-cut partitioners (e.g., hMETIS (KarypisAKS99, ), TritonPart (BustanyGKKPW23, )) that ignore floorplan feasibility, ChipletPart employs a reach-aware, annealing-based placement engine to guarantee floorplan-legal solutions (Section 5.2).
@@
  • Heterogeneous technology assignment: We handle technology node assignment during partitioning using a genetic algorithm (GA). This enables exploration of cost trade-offs in heterogeneous 2.5D systems — a capability not supported by existing tools such as FloorPlet (ChenLZZL23, ) or Chipletizer (Li24, ). Our heterogeneous technology-aware, cost-driven, multi-way partitioner, ChipletPart, finds cost reductions of up to $43\%$ (6% geometric mean) in multi-technology scenarios. To the best of our knowledge, we are the first work that does technology assignment during partitioning Section 3). In addition, we explore the use of Bayesian optimization (BO) to search for floorplan-feasible, cost-optimal chiplet solutions with technology assignments. While BO improves system cost by up to 5.3% over our GA-based framework on a few testcases, it incurs a runtime overhead of up to $4\times$. Thus, GA and BO provide a choice of quality-runtime trade-offs (Section 6.4).
@@
  • Improvements over SOTA: ChipletPart achieves up to 46% improvement in cost over state-of-the-art min-cut based partitioners (hMETIS (KarypisAKS99, ), TritonPart (BustanyGKKPW23, )) that do not guarantee floorplan feasibility. Compared to the parChiplet partitioner from Floorplet (ChenLZZL23, ), ChipletPart generates chiplet solutions that are up to $47\%$ (6% geometric mean) better in cost. Compared to manual partitions, ChipletPart generates $24\%$ (7% geometric mean) better solutions (Section 6). Additionally, we explore how variations in technology parameters affect chiplet partitioning (Section 7).
@@
  • Standardized open-source ecosystem: We translate the Python-based cost model (Graening_CATCH, ) (costmodelUCLA, ) to an equivalent C++ implementation (Section 4.1). This yields a $5\times$ speedup in cost model execution runtime, and a speedup of more than $100\times$ in our implementation. ChipletPart with its core cost model is permissively open-sourced (anonymousrepo, ), enabling others to readily adapt it for benchmarking and further development. Also, we make our testcases publicly available, enabling the community to access a new, standardized set of benchmarks.
@@ 2.1. Chiplet Cost Modeling
Cost reduction is one of the main drivers for developing 2.5D systems. The total cost of developing a VLSI system can be divided into two parts: non-recurring engineering (NRE) cost and recurring engineering (RE) cost (Feng22, ). NRE cost refers to the one-time cost of designing a VLSI system, including IP qualification, architecture, verification, physical design, software, etc. RE cost refers to the fabrication costs in mass production, such as wafers, packaging and testing. Researchers have proposed several chiplet cost models to estimate various components of the total 2.5D system cost. (Feng22, ) introduces a quantitative cost model for comparing RE and NRE costs between monolithic SoC and multi-chip integration, but accounts for fewer RE cost considerations than our chosen model. (ZhuangYCH22, ) and (ChenLZZL23, ) propose cost models that take into account reliability issues such as bump stress and warpage. (GraeningPG23, ; Graening_CATCH, ) presents a case study of a large system built using chiplets and analyzes the sensitivity of system cost to factors such as defect density, assembly cost, IO size, etc. The authors have made their cost model publicly available at (costmodelUCLA, ). In this work, we use the cost model proposed by (costmodelUCLA, ) (see Section 4) due to the wide range of configurable parameters available for system technology co-optimization (STCO). [^2] To the best of our knowledge, the cost model in (costmodelUCLA, ) is the most detailed open-source cost model currently available for chiplets. (Graening_CATCH, ) was used in collaboration with IMEC to perform an STCO study in (Graening24, ). While the specific parameter settings in the cost model do influence partitioning outcomes, all parameters are externally configurable via user-defined configuration files. To enable efficient integration with our partitioning framework, we ported the cost model to C++, resulting in significantly improved performance.
@@
$tot_{pop}$ #members in the population (default 50)
@@
$\sigma$ #members picked for elitism (default 5)
$\Delta_{threshold}$ Threshold value for improvement between
successive generations (default 0.01)
@@ 2.2. Chiplet Partitioning and Floorplanning
Several previous works address chiplet partitioning. (Kabir20, ) incorporates the min-cut partitioner hMETIS (KarypisK96, ) within a chiplet implementation flow. (ChenLZZL23, ) proposes parChiplet, which partitions the SoC system into chiplets based on functional and area characteristics of IP blocks. However, (Kabir20, ) (ChenLZZL23, ) do not optimize the actual chiplets cost. Chipletizer (Li24, ) proposes a unified design characterization graph to represent both SoC designs and chiplets, and uses simulated annealing (SA) to optimize the overall cost of chiplet-based systems. However, Chipletizer does not consider floorplan feasibility or heterogeneous integration. The recent work of (VicharraCH25, ) uses reinforcement learning (RL) and simulated annealing (SA) to perform PPA-oriented chiplet partitioning. However their method does not guarantee floorplan feasibility. The main differences between our ChipletPart and previous works are summarized in Table 3. In our experimental evaluations, we compare ChipletPart with  (KarypisK96, ) and (ChenLZZL23, ). The authors of (Kabir20, )(Li24, ) and (VicharraCH25, ) have not released their source code or binaries. [^3] We attempted to contact the authors of (VicharraCH25, ), but have not received a response.
Table 3. Comparison of chiplet partitioning methods.
Methods
Cost-Driven
Floorplan
Feasibility
Heterogeneous
Integration
@@ Methods
Existing chiplet floorplanning approaches fall into three categories: simulated annealing (SA)-based, branch-and-bound (B&B)-based and mathematical programming (MP)-based. Works such as (HoC13, ) (LeeC23, ) use classical floorplan representations such as B tree and apply SA to optimize an objective function. (ChiouJCLP23, ) (LiuCW14, ) enumerate possible floorplanning solutions and apply B&B to find a near-optimal solution. (ChenLZZL23, ) (ZhuangYCH22, ) develop MP-based formulations of chiplet floorplanning, enabling MP solvers to find optimal solutions. We note that existing works fail to handle the “reach” constraints imposed by IO cells, and generally assume that the size and shape of chiplets are predetermined by designers. Our present work proposes a reach-aware chiplet floorplanner, which simultaneously determines the location and shape for each chiplet.
@@ 3.1. Problem Formulation
Chiplet partitioning is fundamentally different from classical min-cut partitioning. The key differences are summarized as follows.
  • Problem size: Min-cut partitioning is typically performed on gate-level netlists comprising millions of gates. By contrast, chiplet partitioning focuses on a block-level netlist that typically comprises several hundreds of IP blocks.
  • Objective: Min-cut partitioning focuses on minimizing cutsize, the number of nets (hyperedges) crossing partition boundaries. By contrast, chiplet partitioning seeks to reduce 2.5D system development cost, which brings a more complex set of considerations (Section  4).
  • Constraints: Min-cut partitioning minimizes cutsize subject to a given balance constraint (BustanyGKKPW23, ). By contrast, chiplet partitioning is not necessarily balanced, but it must produce IO-feasible solutions (Section 5.2).
@@
where $\phi_{die}$ is the silicon cost of each chiplet, $\phi_{assembly}$ is the assembly cost, $Y_{die}$ is the die yield, $Y_{assembly}$ is the assembly yield, $g(\omega)$ is the non-recurring engineering cost, and $V$ is the chip manufacturing volume. The constraint is the IO feasibility of the chiplet set $\mathcal{C}$.
@@ 3.2. Overview of ChipletPart Framework
We now describe our chiplet partitioning framework, ChipletPart (see Figure 2). Our ChipletPart differs from classical min-cut partitioners (KarypisAKS99, ; bustany2022specpart, ; BustanyGKKPW23, ) and previous approaches (Sangiovanni-VincentelliLZZ23, ; Li24, ; ChenLZZL23, ; Kabir20, ) in several key respects. (i) We optimize a comprehensive chiplet cost function (Section 4), unlike min-cut partitioners that focus on minimizing net cutsize. (ii) We integrate a fast, Go-With-the-Winners (Aldous94, ) SA-based chiplet floorplanner to ensure IO-feasibility of each partitioning solution (Section 5.2). (iii) We leverage a genetic algorithm (GA) framework to discover high-quality technology node assignments for heterogeneous integration.
@@
Step 1: [Lines 3-4] We first generate the initial population in the GA. Each member in the population (a genome) represents an ordered sequence of technology assignments (to chiplets) $\omega:\mathcal{C}\to\mathcal{T}$. We generate $tot_{pop}$ ($tot_{pop}$ = 50 by default) genomes; each genome is generated by randomly mapping $K_{\text{max}}$ chiplets to technology nodes $\mathcal{T}$}. [^5] For simplicity, we set $tot_{pop}=k_{pop}+\sigma$. To avoid redundant evaluations and improve scalability, we canonicalize all generated genomes by mapping functionally equivalent assignments to a unique representation. For example, genomes $\langle 7nm,7nm,14nm\rangle$ and $\langle 14nm,7nm,7nm\rangle$ are treated as equivalent, since they correspond to the same multiset of assignments.
Figure 3. GA-based technology assignment. A genome is a set of technology nodes. Shown: two parent genomes $\langle 7,7,14\rangle$ and $\langle 10,10,10\rangle$ undergo crossover to produce the offspring $\langle 7,10,14\rangle$; the offspring undergoes mutation to produce $\langle 10,10,7\rangle$ for the next-generation population.
[…]
Step 2: [Lines 5-11] We assess the fitness of each genome in the current population. For each genome $\omega_{j}$, we (i) run Core-ChipletPart [^6] For better scalability, we run Core-ChipletPart with fewer FM (Fiduccia–Mattheyses) moves and passes, and fewer initial solutions. Details are seen in our code (anonymousrepo, ). to generate a partitioning solution based on $\omega_{j}$ and (ii) calculate the fitness score of this solution using our cost model.
@@
Step 5: [Lines 23-25] We generate the next-generation population (offsprings) through crossover, elitism, and mutation. Notably, we use uniform crossover (syswerda1989uniform, ) and random resetting mutation  (deb1999introduction, ). Additionally, we retain the top $\sigma$ genomes from the parent generation based on their fitness scores (elitism), ensuring that high-quality solutions persist across generations.
In our implementation, we set $tot_{pop}=50$, $\zeta=3$, $k_{pop}=45$, $\sigma=5$ and $K_{max}=8$ (Section 6.2.4). So, we conduct $90$ tournaments (each tournament picks 3 genomes from the current generation) to select $90$ parent genomes, which are then paired to generate $45$ offspring genomes through crossover and mutation. We then supplement these $45$ offspring with the top $5$ genomes from the current generation (elitism), to generate the next generation population of $50$ genomes. For additional implementation details, we refer the reader to (anonymousrepo, ).
Choice of optimizers: Our choice of picking GA and SA is motivated by their ability to flexibly optimize over black-box, non-convex and highly discontinuous cost landscapes — characteristics inherent to our chiplet cost model (Section 4.1). In contrast, ILP-based methods are less suitable for our framework due to the lack of closed-form or linearizable expressions for IO feasibility and chiplet cost functions. GA enables direct co-optimization of technology assignment and chiplet partitioning, while SA ensures that floorplanning solutions are always IO-feasible. Having a unified optimization approach with a unified formulation is challenging since the solution space spans both discrete combinatorial variables (chiplet-to-tech assignment) and continuous geometric constraints (IO reach), which are difficult to encode and optimize jointly. We hence opt for a modular decomposition — using GA for assignment and SA for feasibility evaluation. Bayesian optimization (BO) is a powerful tool for optimizing complicated, black-box functions — we explore its potential as an alternative optimizer in Section 6.4.

### v3

@@ Abstract.
2.5D, Chiplets, Partitioning, Floorplanning, Chiplet Manufacturing Cost Model
journal: TODAESccs: Hardware 3D integrated circuitsccs: Hardware Partitioning and floorplanningccs: Hardware Economics of chip design and manufacturing
@@ 1. Introduction
The integration of multiple chips on an interposer has become a favorable approach to reduce the cost of building large systems (IntelPV, ; Sangiovanni-VincentelliLZZ23, ). In a 2.5D system, a design is decomposed into multiple smaller chiplets, which are packaged together on the same substrate. This splitting can have significant benefits for yield and can enable designs using heterogeneous process technology nodes, going beyond what is possible in a monolithic design. On the other hand, inter-die communication exhibits higher area and power overhead compared to intra-die communication. Consequently, the potential cost benefits of disaggregation are not always realized (GraeningPG23, ). Therefore, it is important to intelligently partition a design into constituent chiplets and assign a manufacturing technology to each chiplet so as to optimize manufacturing costs.
@@
The natural partitioning granularity for 2.5D integration is at the block-level. Splitting individual IP blocks into multiple chiplets is problematic for design and test methodology reasons as well as for IP reuse. Additionally, splitting an IP block will often have a significant impact on performance and I/O count. Due to these considerations, block-level chiplet partitioning (Figure 1) is the focus of this paper. The cost-aware partitioning problem for chiplet systems is fundamentally different from the well-studied netlist min-cut partitioning (KarypisAKS99, ; bustany2022specpart, ; BustanyGKKPW23, ). For the former, the problem size is smaller (usually a few hundred blocks and a few tens of chiplets at most) and the underlying objective function can be complex. Furthermore, inter-chiplet I/O reach limitations can render some partitions infeasible, necessitating floorplan-awareness in the partitioner. This is further complicated by the need to select a manufacturing technology for each chiplet.
Refer to caption An image showing colored circles on the left representing blocks being partitioned into colored rectangles representing chiplets on the right.
@@
Figure 1. Generic example of block-level chiplet partitioning. Left: an IP block-level netlist (different types of IP blocks shown in different colors). Right: an integrated 2.5D chiplet system reflecting the block-level netlist partitioning with technology node assignment shown in different colors.
@@
We propose ChipletPart, the first open-source, unified framework for chiplet partitioning that combines cost-driven multi-way partitioning, heterogeneous technology assignment and I/O reach-aware floorplanning in a single, automated flow. ChipletPart presents a novel adaptation of classical optimizers such as genetic algorithms (GA) and simulated annealing (SA) to generate floorplan-feasible, cost-optimized 2.5D chiplet system partitions with technology assignments. ChipletPart is intended for design space exploration before physical implementation. [^1] ChipletPart considers a given architecture and generates partitioned chiplets before physical implementation. We do not perform architecture exploration in this work. We envision ChipletPart as a tool to assist human designers by generating strong initial solutions, which can then be refined using domain expertise. Our key contributions include:
@@
  • Integrated chiplet partitioning with floorplan-feasibility: We present the first unified framework that optimizes system partitioning while guaranteeing feasibility under I/O reach constraints. In contrast to standard min-cut partitioners (e.g., hMETIS (KarypisAKS99, ), TritonPart (BustanyGKKPW23, )) that ignore floorplan-feasibility, ChipletPart employs a reach-aware, simulated annealing-based placement engine to guarantee floorplan-feasible solutions (Section 5.2).
@@
  • Heterogeneous technology assignment: We handle technology node assignment during partitioning using a genetic algorithm. This enables exploration of cost trade-offs in heterogeneous 2.5D systems — a capability not supported by existing tools such as Floorplet (ChenLZZL23, ) or Chipletizer (Li24, ). Our heterogeneous technology-aware, cost-driven, multi-way partitioner, ChipletPart, finds cost reductions of up to $43\%$ (6% geometric mean) in multi-technology scenarios. To the best of our knowledge, we are the first work that does technology assignment during partitioning (Section 3). In addition, we explore the use of Bayesian optimization (BO) to search for floorplan-feasible, cost-optimal chiplet solutions with technology assignments. While BO improves system cost by up to 5.3% over our GA-based framework on a few testcases, it incurs a runtime overhead of up to $4\times$. Thus, GA and BO provide a quality-runtime trade-off (Section 6.4).
@@
  • Improvements over SOTA: ChipletPart achieves up to 46% improvement in cost over state-of-the-art min-cut based partitioners (hMETIS (KarypisAKS99, ), TritonPart (BustanyGKKPW23, )) that do not guarantee floorplan-feasibility. Compared to the parChiplet partitioner from Floorplet (ChenLZZL23, ), ChipletPart generates chiplet solutions that are up to $47\%$ (6% geometric mean) better in cost. Compared to manual partitions, ChipletPart generates $34\%$ (13% geometric mean) better solutions (Section 6). We also compare against the cost-driven Chipletizer (Li24, ) framework and observe that ChipletPart consistently produces I/O-feasible partitions with up to 48% lower cost across our design suite (Section 6). Additionally, we explore how variations in technology parameters affect chiplet partitioning (Section 7).
@@
  • Standardized open-source ecosystem: We translated the Python-based cost model (Graening_CATCH, ; costmodelUCLA, ) to an equivalent C++ implementation (Section 4.1). This yields a $5\times$ speedup in cost model execution runtime, and a speedup of more than $100\times$ in our multithreaded implementation. ChipletPart, along with its core cost model, is permissively open-sourced (anonymousrepo, ), enabling others to readily adapt it for benchmarking and further development. Also, we make our testcases publicly available, enabling the community to access a new, standardized set of benchmarks.
@@ 2.1. Chiplet Cost Modeling
Cost reduction is one of the main drivers for developing 2.5D systems. The total cost of developing a VLSI system can be divided into two parts: non-recurring engineering (NRE) cost and recurring engineering (RE) cost (Feng22, ). NRE cost refers to the one-time cost of designing a VLSI system, including IP qualification, architecture simulation, verification, physical design, software license fees, etc. RE cost refers to the fabrication costs in mass production, such as wafers, assembly, and test. Researchers have proposed several chiplet cost models to estimate various components of the total 2.5D system cost. (Feng22, ) introduces a quantitative cost model for comparing RE and NRE costs between monolithic SoC and multi-chip integration, but accounts for fewer RE cost considerations than our chosen model. (ZhuangYCH22, ) and (ChenLZZL23, ) propose cost models that take into account reliability issues such as bump stress and warpage. (GraeningPG23, ; Graening_CATCH, ) present a case study of a large system built using chiplets and analyze the sensitivity of system cost to factors such as defect density, assembly cost, I/O size, etc. The authors have made their cost model publicly available at (costmodelUCLA, ). In this work, we use the cost model proposed by (costmodelUCLA, ) (see Section 4) due to the wide range of configurable parameters available for system technology co-optimization (STCO). [^2] To the best of our knowledge, the cost model in (costmodelUCLA, ) is the most detailed open-source cost model currently available for chiplets at the time of this publication. The cost model in (Graening_CATCH, ) was used in collaboration with IMEC to perform an STCO study in (Graening24, ). While the specific parameter settings in the cost model do influence partitioning outcomes, all parameters are externally configurable via user-defined configuration files. To enable efficient integration with our partitioning framework, we ported the cost model to C++, resulting in significantly improved performance.
@@
$tot_{pop}$ Number of members in the population (default 50)
@@
$\sigma$ Number of members picked for elitism (default 5)
$\Delta_{threshold}$ Threshold value for improvement between successive generations (default 0.01)
@@ 2.2. Chiplet Partitioning and Floorplanning
Several previous works address chiplet partitioning. (Kabir20, ) incorporates the min-cut partitioner hMETIS (KarypisK96, ) within a chiplet implementation flow. (ChenLZZL23, ) proposes parChiplet, which partitions the SoC system into chiplets based on functional and area characteristics of IP blocks. However, (Kabir20, ; ChenLZZL23, ) do not optimize the actual chiplet cost. Chipletizer (Li24, ) proposes a unified design characterization graph to represent both SoC designs and chiplets, and uses simulated annealing (SA) to optimize the overall cost of chiplet-based systems. However, Chipletizer does not consider floorplan feasibility or support heterogeneous integration. The recent work of (VicharraCH25, ) uses reinforcement learning (RL) and simulated annealing to perform PPA-oriented chiplet partitioning. However, their method does not guarantee floorplan-feasibility. The main differences between our ChipletPart and previous works are summarized in Table 3. In our experimental evaluations, we compare ChipletPart with  (ChenLZZL23, ) and (KarypisK96, ). The authors of (Li24, ; Kabir20, ; VicharraCH25, ) have not released their source code or binaries. [^3] We attempted to contact the authors of (VicharraCH25, ), but have not received a response.
Table 3. Comparison of state-of-the-art chiplet partitioning methods.
Methods Cost-Driven Floorplan Feasibility Heterogeneous Integration
@@
Existing chiplet floorplanning approaches fall into three categories: simulated annealing-based, branch-and-bound (B&B)-based and mathematical programming (MP)-based. Works such as (HoC13, ; LeeC23, ) use classical floorplan representations such as B tree and apply SA to optimize an objective function. (ChiouJCLP23, ; LiuCW14, ) enumerate possible floorplanning solutions and apply B&B to find a near-optimal solution. (ChenLZZL23, ; ZhuangYCH22, ) develop MP-based formulations of chiplet floorplanning, enabling MP solvers to find optimal solutions. We note that existing works fail to handle the “reach” constraints imposed by I/O cells, and generally assume that the size and shape of chiplets are predetermined by designers. In this work, we propose a reach-aware chiplet floorplanner that simultaneously determines the location and shape for each chiplet.
@@ 3.1. Problem Formulation
The chiplet partitioning problem that we address is fundamentally different from classical min-cut partitioning. The key differences are summarized as follows.
  • Problem size: Min-cut partitioning is typically performed on gate-level netlists comprising millions of gates. By contrast, we focus on a block-level netlist that typically comprises several hundreds of IP blocks.
  • Objective: Min-cut partitioning focuses on minimizing cutsize, the number of nets (hyperedges) crossing partition boundaries. By contrast, we seek to reduce 2.5D system development cost, which brings a more complex set of considerations (Section  4).
  • Constraints: Min-cut partitioning minimizes cutsize subject to a given balance constraint (BustanyGKKPW23, ). By contrast, chiplet partitioning is not necessarily balanced, but it must produce I/O-feasible solutions (Section 5.2).
@@
where $\phi_{die}$ is the silicon cost of each chiplet, $\phi_{assembly}$ is the assembly cost, $Y_{die}$ is the die yield, $Y_{assembly}$ is the assembly yield, $g(\omega)$ is the non-recurring engineering (NRE) cost, and $V$ is the chip manufacturing volume. The constraint is the I/O-feasibility of the chiplet set $\mathcal{C}$.
Refer to caption A flowchart diagram showing the inputs and outputs of the ChipletPart framework.
@@ 3.2. Overview of ChipletPart Framework
We now describe our chiplet partitioning framework, ChipletPart (see Figure 2). Our ChipletPart differs from classical min-cut partitioners (KarypisAKS99, ; bustany2022specpart, ; BustanyGKKPW23, ) and previous approaches (Sangiovanni-VincentelliLZZ23, ; Li24, ; ChenLZZL23, ; Kabir20, ) in several key respects. (i) We optimize a comprehensive chiplet cost function (Section 4), unlike min-cut partitioners that focus on minimizing net cutsize. (ii) We integrate a fast, Go-With-the-Winners (Aldous94, ) SA-based chiplet floorplanner to ensure I/O-feasibility of a partitioning solution (Section 5.2). (iii) We leverage a genetic algorithm (GA) framework to discover high-quality technology node assignments for heterogeneous integration.
@@
Step 1: [Lines 3-4] We first generate the initial population in the GA. Each member in the population (a genome) represents an ordered sequence of technology assignments (to chiplets) $\omega:\mathcal{C}\to\mathcal{T}$. We generate $tot_{pop}$ ($tot_{pop}$ = 50 by default) genomes; each genome is generated by randomly mapping $K_{\text{max}}$ chiplets to technology nodes $\mathcal{T}$}. [^5] For simplicity, we set $tot_{pop}=k_{pop}+\sigma$. To avoid redundant evaluations and improve scalability, we canonicalize all generated genomes by mapping functionally equivalent assignments to a unique representation. For example, genomes $\langle 7\,\text{nm},7\,\text{nm},14\,\text{nm}\rangle$ and $\langle 14\,\text{nm},7\,\text{nm},7\,\text{nm}\rangle$ are treated as equivalent, since they correspond to the same multiset of assignments.
Refer to caption A diagram showing the behavior of the genetic algorithm including crossover and mutation.
Figure 3. GA-based technology assignment. A genome is a set of technology nodes. Shown: two parent genomes $\langle 7,7,14\rangle$ and $\langle 10,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}10},10\rangle$ undergo crossover to produce the offspring $\langle 7,10,14\rangle$; the offspring undergoes mutation to produce $\langle 10,10,7\rangle$ for the next-generation population.
Step 2: [Lines 5-11] We assess the fitness of each genome in the current population. For each genome $\omega_{j}$, we (i) run Core-ChipletPart [^6] For better scalability, we run Core-ChipletPart with fewer FM (Fiduccia-Mattheyses) moves and passes, and fewer initial solutions. Details are seen in our code (anonymousrepo, ). to generate a partitioning solution based on $\omega_{j}$ and (ii) calculate the fitness score of this solution using our cost model.
@@
Step 5: [Lines 23-25] We generate the next-generation population (offspring) through crossover, elitism, and mutation. Notably, we use uniform crossover (syswerda1989uniform, ) and random resetting mutation  (mitchell1998introduction, ). Additionally, we retain the top $\sigma$ genomes from the parent generation based on their fitness scores (elitism), ensuring that high-quality solutions persist across generations.
In our implementation, we set $tot_{pop}=50$, $\zeta=3$, $k_{pop}=45$, $\sigma=5$ and $K_{max}=8$ (Section 6.2.5). So, we conduct $90$ tournaments (each tournament picks three genomes from the current generation) to select $90$ parent genomes, which are then paired to generate $45$ offspring genomes through crossover and mutation. We then supplement these $45$ offspring with the top five genomes from the current generation (elitism) to generate the next generation population of $50$ genomes. For additional implementation details, we refer the reader to (anonymousrepo, ).
Choice of optimizers: Our choice of picking GA and SA is motivated by their ability to flexibly optimize over black-box, non-convex and highly discontinuous cost landscapes — characteristics inherent to our chiplet cost model (Section 4.1). In contrast, ILP-based methods are less suitable for our framework due to the lack of closed-form or linearizable expressions for I/O-feasibility and chiplet cost functions. GA enables direct co-optimization of technology assignment and chiplet partitioning, while SA ensures that floorplanning solutions are always I/O-feasible. Having a unified optimization approach and formulation is challenging since the solution space spans both discrete combinatorial variables (chiplet-to-tech assignment) and continuous geometric constraints (I/O reach), which are difficult to encode and optimize jointly. We hence opt for a modular decomposition — using GA for assignment and SA for feasibility evaluation. Note that Bayesian optimization (BO) is another powerful tool for optimizing complicated, black-box functions — we explore its potential as an alternative optimizer in Section 6.4.